OpenStudy (anonymous):

Find all solutions of the equation in the interval [0°, 360°). cos(x)=-sq route of 2 over 2

OpenStudy (solomonzelman):

Okay, Cos(45)= S / S√2 = 1/√2 = √2/2 SO cosine of what is =√2/2 ?